# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ 3 A single-loop SPIDER-type stochastic subgradient algorithm
In this section, we present our algorithm for solving problem Eq. P. It is developed based on a subgradient method for problem Eq. PP by using the SPIDER technique to estimate constraint function values. The algorithm framework is outlined in Section 3.1. In Section 3.2, we prove that in expectation, the algorithm can find a near-$\epsilon$ stationary point of problem Eq. PP. Building on the results presented in the previous section, we further establish in Section 3.3 that an $(\epsilon,\epsilon)$-KKT point in expectation of problem Eq. P can be obtained.
@@ 3.1 Algorithm framework
We employ the SPIDER technique [16] to obtain an estimator $u^{(k)}$ of $g$ at the $k$-th iteration. Specifically, every $q$ iterations, we draw a big batch of $S_{1}$ samples from $\mathbb{P}_{g}$, while for other iterations, a smaller batch of $S_{2}$ samples is drawn. The estimator is set as
(25) $u^{(k)}=\left\{\begin{array}[]{lll}g({\mathbf{x}}^{(k)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{1},&\text{ if }\mathrm{mod}(k,q)=0,\\[2.84526pt] u^{(k-1)}+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{2},&\text{ otherwise},\end{array}\right.$
where $\mathcal{B}_{k}$ contains identically independent samples from $\mathbb{P}_{g}$, and $g(\mathbf{x},\mathcal{B})$ is defined in Eq. 6. With $u^{(k)}$, we then take a sample subgradient of $f$ and $g$ at ${\mathbf{x}}^{(k)}$ and perform the update in Eq. 1. The pseudocode is given in Algorithm 1.
Algorithm 1 A SPIDER-type Stochastic Subgradient algorithm
1:  Input: initial point ${\mathbf{x}}^{(0)}\in{\mathrm{dom}}(f)\cap{\mathrm{dom}}(g)$, $\beta>0$, $q>0$, $S_{1},S_{2}>0$, and $K>0$.
@@
3:     Set $u^{(k)}$ by Eq. 25.
4:     Choose $\alpha_{k}>0$, generate a (stochastic) subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and a (stochastic) subgradient $\boldsymbol{\zeta}_{g}^{(k)}$ of $g$.
@@
(26) $\boldsymbol{\zeta}_{F^{\nu}}^{(k)}=\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\text{, and }$
${\mathbf{w}}^{(k)}=\left(\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g}^{(k)}.$
@@
(27) ${\mathbf{x}}^{(k+1)}={\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right).$
Throughout this section, we make the following assumption, where for brevity of notations, we let $\mathbb{E}_{k}[\cdot]:=\mathbb{E}\left[\cdot\mid{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\right]$, representing the conditional expectation given $\{{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\}$ for any $k\geq 0$. In addition, $\mathrm{Var}(\boldsymbol{\zeta}):=\mathbb{E}[\|\boldsymbol{\zeta}-\mathbb{E}[\boldsymbol{\zeta}]\|^{2}]$ for a random vector $\boldsymbol{\zeta}$.
@@
    Unbiased function estimate and Lipschitz continuity of $g$: For all $k\geq 0$, the samples in $\mathcal{B}_{k}$ are mutually independent. It holds $\mathbb{E}_{k}[g({\mathbf{x}}^{(k)},\xi_{g})]=g({\mathbf{x}}^{(k)})$ and $\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\sigma^{2}$ for any $\xi_{g}\in\mathcal{B}_{k}$ and some $\sigma>0$. Moreover, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathrm{dom}}(g)$, it holds $\mathbb{E}\left[\left|g({\mathbf{x}},\xi_{g})-g({\mathbf{y}},\xi_{g})\right|^{2}\right]\leq L_{g}^{2}\|{\mathbf{x}}-{\mathbf{y}}\|^{2}.$
@@
    Unbiased Stochastic Subgradients: For each $k$, $\mathbb{E}_{k}[\boldsymbol{\zeta}_{f}^{(k)}]\in\partial f({\mathbf{x}}^{(k)})$ and $\mathbb{E}_{k}[\boldsymbol{\zeta}_{g}^{(k)}]\in\partial g({\mathbf{x}}^{(k)})$. In addition, there are $\sigma_{1}\geq 0$ and $\sigma_{2}\geq 0$ such that $\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})\leq\sigma_{1}^{2}$ and $\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})\leq\sigma_{2}^{2}$.
By 4(b), it follows that $\mathbb{E}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$.
3.2 Convergence analysis and complexity results
In this subsection, we present the convergence analysis of Algorithm 1. First, we derive a bound on the difference between $u^{(k)}$ and $g({\mathbf{x}}^{(k)})$.
@@ 3.2 Convergence analysis and complexity results
Under 1 and 4, let $\{{\mathbf{x}}^{(k)},u^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1. Define $r_{g}^{(k)}:=\mathbb{E}\left[|u^{(k)}-g({\mathbf{x}}^{(k)})|^{2}\right]$. Then, for all $k\geq 0$, it holds $r_{g}^{(k)}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
@@
Case (i): When $\mathrm{mod}(k,q)=0$,  Eq. 25 and 4(a) give
(28) $r_{g}^{(k)}=\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]\leq\frac{\sigma^{2}}{S_{1}}.$
@@
$=r_{g}^{(k-1)}+\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k-1)},\mathcal{B}_{k})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
$=r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k-1)})+g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})-g({\mathbf{x}}^{(k)})\right|^{2}\right]$
$\leq r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{g}\in\mathcal{B}_{k}}\mathbb{E}\left[\left|g({\mathbf{x}}^{(k)},\xi_{g})-g({\mathbf{x}}^{(k-1)},\xi_{g})\right|^{2}\right]$
$\leq r_{g}^{(k-1)}+\frac{L_{g}^{2}}{S_{2}}\mathbb{E}\left[\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{(k-1)}\|^{2}\right]\leq r_{g}^{(k-1)}+\frac{\alpha_{k-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}},$
where the second equality holds due to the unbiasedness of $g({\mathbf{x}},\xi_{g})$, the third equality follows from the mutual independence of samples in $\mathcal{B}_{k}$, the second inequality follows by 4(a), and the last one uses
(30) $\mathbb{E}\left[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|^{2}\right]=\textstyle\alpha_{k}^{2}\mathbb{E}\left[\left\|\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\mathrm{Proj}_{[0,1]}\left(\frac{u^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g}^{(k)}\right)\right\|^{2}\right]$
$\leq 2\alpha_{k}^{2}\left(\mathbb{E}\big{[}\|\boldsymbol{\zeta}_{f}^{(k)}\|^{2}+\beta^{2}\|\boldsymbol{\zeta}_{g}^{(k)}\|^{2}\big{]}\right)$
$\leq 2\alpha_{k}^{2}\left(\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})+\|\mathbb{E}[\boldsymbol{\zeta}_{f}^{(k)}]\|^{2}+\beta^{2}\mathrm{Var}(\boldsymbol{\zeta}_{g}^{(k)})+\beta^{2}\|\mathbb{E}[\boldsymbol{\zeta}_{g}^{(k)}]\|^{2}\right)$
$\leq 2\alpha_{k}^{2}(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})=\alpha_{k}^{2}L_{F}^{2}.$
Since $\mathrm{mod}(k,q)>0$, there exist nonnegative integers $t$ and $s<q$ such that $k=tq+s$. Summing Eq. 29 for all $k=tq+1,\ldots,tq+s-1,tq+s$, and applying Eq. 28, we have
@@
The proof is then completed by combining the above two cases.
The following corollary is obtained immediately from Eq. 31 by noticing $r_{g}^{(tq)}=0$ if $\mathbb{P}_{g}$ is a uniform distribution over $N$ data points and ${\mathcal{B}}_{tq}$ contains all the data points.
@@
Under 1 and 4, suppose that $\mathbb{P}_{g}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Let ${\mathcal{B}}_{k}=\Xi$ if $\mathrm{mod}(k,q)=0$. Then $r_{g}^{(k)}=0$ if $\mathrm{mod}(k,q)=0$ and $r_{g}^{(k)}\leq\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$ otherwise, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$.
Next, we bound the difference between the Moreau envelope at two consecutive iterates.
@@
Under 1 and 4, suppose that $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{u^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1. With $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}\sigma_{2}^{2}+\beta^{2}l_{g}^{2})$, it holds that
@@
where the first inequality follows directly from the definition of $F^{\nu}_{\frac{1}{2\rho}}$, and the last one uses $\mathbb{E}_{k}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$ by 4(b) and the $\rho$-weak convexity of $F^{\nu}$ by Lemma 2.6.

### v2

@@ 3 A SPIDER-type stochastic subgradient algorithm
In this section, we present our algorithm, termed 3S-Econ, for solving problem Eq. P. It is developed based on a subgradient method for problem Eq. PP by using the SPIDER technique to estimate constraint function values. The algorithm framework is outlined in Section 3.1. In Section 3.2, we prove that in expectation, 3S-Econ can find a near-$\epsilon$ stationary point of problem Eq. PP. Building on the results presented in the previous section, we further establish in Section 3.3 that an $(\epsilon,\epsilon)$-KKT point in expectation of problem Eq. P can be obtained.
@@ 3.1 Algorithm framework
We employ the SPIDER technique [24] to obtain an estimator ${\mathbf{u}}^{(k)}\in\mathbb{R}^{m}$ of $\mathbf{g}$} at the $k$-th iteration. Specifically, every $q$ iterations, we draw a big batch of $S_{1}$ samples from $\mathbb{P}_{{\mathbf{g}}}$, while for other iterations, a smaller batch of $S_{2}$ samples is drawn. The estimator is set as
(26) ${\mathbf{u}}^{(k)}=\left\{\begin{array}[]{lll}{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{1},&\text{ if }\mathrm{mod}(k,q)=0,\\[2.84526pt] {\mathbf{u}}^{(k-1)}+{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\mathcal{B}_{k}),&\text{ with }|\mathcal{B}_{k}|=S_{2},&\text{ otherwise},\end{array}\right$.}
where ${\mathbf{u}}^{(k)}=[u_{1}^{(k)},\ldots,u_{m}^{(k)}]^{\top}$, $\mathcal{B}_{k}$ contains identically independent samples from $\mathbb{P}_{{\mathbf{g}}}$, and $g_{i}(\mathbf{x},\mathcal{B})$ is defined in Eq. 7. With ${\mathbf{u}}^{(k)}$, we then take a sample subgradient of $f$ and $\{g_{i}\}_{i\in[m]}$ at ${\mathbf{x}}^{(k)}$ and perform the update in Eq. 2. The pseudocode is given in Algorithm 1. As detailed in Theorem 3.6, for a given error tolerance $\bar{\epsilon}>0$, we set the penalty parameter $\beta$ to a sufficiently large predefined constant, the smoothing parameter $\nu=O(\bar{\epsilon})$, the step size $\alpha_{k}=O(\bar{\epsilon}^{2})$, and $q=S_{2}=\lceil\sqrt{S_{1}}\rceil$, with $S_{1}=O(\bar{\epsilon}^{-4})$ in the generic stochastic case or $S_{1}=N$ for the finite-sum case.
Algorithm 1 3S-Econ: A SPIDER-type Stochastic Subgradient algorithm for solving Expectation constrained problem (P)
1:Input: ${\mathbf{x}}^{(0)}\in\mathcal{X}$, $\beta>0$, $\nu>0$, and positive intergers $q$, $S_{1},S_{2}$, and $K$.
@@
3:  Sample ${\mathcal{B}}_{k}$ and set ${\mathbf{u}}^{(k)}$ by Eq. 26.
4:  Choose $\alpha_{k}>0$, generate a (stochastic) subgradient $\boldsymbol{\zeta}_{f}^{(k)}$ of $f$ and (stochastic) subgradient $\boldsymbol{\zeta}_{g_{i}}^{(k)}$ of $g_{i}$ for each $i\in[m]$ that are independent of ${\mathcal{B}}_{k}$.
@@
(27) $\boldsymbol{\zeta}_{F^{\nu}}^{(k)}=\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\text{, and }$
${\mathbf{w}}^{(k)}=\sum_{i=1}^{m}\left(\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)-\mathrm{Proj}_{[0,1]}\left(\frac{g_{i}({\mathbf{x}}^{(k)})}{\nu}\right)\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}.$
@@
(28) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathbf{x}}^{(k+1)}=\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right).$
Throughout this section, we make the following assumption, where for brevity of notation, we let $\mathbb{E}_{k}[\cdot]:=\mathbb{E}\left[\cdot\mid{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\right]$, representing the conditional expectation given $\{{\mathbf{x}}^{(0)},{\mathbf{x}}^{(1)},\ldots,{\mathbf{x}}^{(k)}\}$ for any $k\geq 0$. In addition, $\mathrm{Var}(\boldsymbol{\zeta}):=\mathbb{E}[\|\boldsymbol{\zeta}-\mathbb{E}[\boldsymbol{\zeta}]\|^{2}]$ for a random vector $\boldsymbol{\zeta}$.
@@
    Unbiased function estimate and Lipschitz continuity of $\mathbf{g}$}: For all $k\geq 0$, the samples in $\mathcal{B}_{k}$ are mutually independent. It holds $\mathbb{E}_{k}[{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})]={\mathbf{g}}({\mathbf{x}}^{(k)})$ and $\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\sigma^{2}$ for any $\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}$ and some $\sigma>0$. Moreover, for all ${\mathbf{x}},{\mathbf{y}}\in{\mathcal{X}}$, it holds $\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{y}},\xi_{{\mathbf{g}}})\right\|^{2}\right]\leq L_{g}^{2}\|{\mathbf{x}}-{\mathbf{y}}\|^{2}.$
@@
    Unbiased Stochastic Subgradients: For each $k$, $\mathbb{E}_{k}[\boldsymbol{\zeta}_{f}^{(k)}]\in\partial f({\mathbf{x}}^{(k)})$ and $\mathbb{E}_{k}[\boldsymbol{\zeta}_{g_{i}}^{(k)}]\in\partial g_{i}({\mathbf{x}}^{(k)})$ for each $i\in[m]$. In addition, there are $\sigma_{1}\geq 0$ and $\sigma_{2}\geq 0$ such that $\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})\leq\sigma_{1}^{2}$ and $\mathrm{Var}(\boldsymbol{\zeta}_{g_{i}}^{(k)})\leq\sigma_{2}^{2}$ for each $i\in[m]$.
By 3(b), it follows that $\mathbb{E}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$.
3.2 Convergence analysis
In this subsection, we present the convergence analysis of Algorithm 1. First, we derive a bound on the difference between ${\mathbf{u}}^{(k)}$ and ${\mathbf{g}}({\mathbf{x}}^{(k)})$.
@@ 3.2 Convergence analysis
Under 1 and 3, let $\{{\mathbf{x}}^{(k)},{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ be generated by Algorithm 1. Define $r_{g}^{(k)}:=\mathbb{E}\left[\|{\mathbf{u}}^{(k)}-{\mathbf{g}}({\mathbf{x}}^{(k)})\|^{2}\right]$. Then, for all $k\geq 0$, it holds $r_{g}^{(k)}\leq\frac{\sigma^{2}}{S_{1}}+\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$, where $L_{F}^{2}=2(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}m^{2}\sigma_{2}^{2}+\beta^{2}m^{2}l_{g}^{2})$.
@@
Case (i): When $\mathrm{mod}(k,q)=0$,  Eq. 26 and 3(a) give
(29) $r_{g}^{(k)}=\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]\leq\frac{\sigma^{2}}{S_{1}}.$
@@
$=r_{g}^{(k-1)}+\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k-1)})+{\mathbf{g}}({\mathbf{x}}^{(k)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\mathcal{B}_{k})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
$=r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}}\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k-1)})+{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k)})\right\|^{2}\right]$
$\leq r_{g}^{(k-1)}+\frac{1}{|\mathcal{B}_{k}|^{2}}\sum_{\xi_{{\mathbf{g}}}\in\mathcal{B}_{k}}\mathbb{E}\left[\left\|{\mathbf{g}}({\mathbf{x}}^{(k)},\xi_{{\mathbf{g}}})-{\mathbf{g}}({\mathbf{x}}^{(k-1)},\xi_{{\mathbf{g}}})\right\|^{2}\right]$
(30) $\leq r_{g}^{(k-1)}+\frac{L_{g}^{2}}{S_{2}}\mathbb{E}\left[\|{\mathbf{x}}^{(k)}-{\mathbf{x}}^{(k-1)}\|^{2}\right]\leq r_{g}^{(k-1)}+\frac{\alpha_{k-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}},$
where the second equality holds due to the unbiasedness of ${\mathbf{g}}({\mathbf{x}},\xi_{{\mathbf{g}}})$, the third equality follows from the mutual independence of samples in $\mathcal{B}_{k}$, the second inequality follows by 3(a), and the last one uses
(31) $\mathbb{E}\left[\|{\mathbf{x}}^{(k+1)}-{\mathbf{x}}^{(k)}\|^{2}\right]=\mathbb{E}\left[\left\|\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right)-\mathrm{Proj}_{\mathcal{X}}\left({\mathbf{x}}^{(k)}\right)\right\|^{2}\right]$
$\leq\textstyle\alpha_{k}^{2}\mathbb{E}\left[\left\|\left(\boldsymbol{\zeta}_{f}^{(k)}+\beta\sum_{i=1}^{m}\mathrm{Proj}_{[0,1]}\left(\frac{u_{i}^{(k)}}{\nu}\right)\boldsymbol{\zeta}_{g_{i}}^{(k)}\right)\right\|^{2}\right]$
$\leq 2\alpha_{k}^{2}\left(\mathbb{E}\left[\|\boldsymbol{\zeta}_{f}^{(k)}\|^{2}+\beta^{2}m\sum_{i=1}^{m}\left\|\boldsymbol{\zeta}_{g_{i}}^{(k)}\right\|^{2}\right]\right)$
$\leq 2\alpha_{k}^{2}\left(\mathrm{Var}(\boldsymbol{\zeta}_{f}^{(k)})+\|\mathbb{E}[\boldsymbol{\zeta}_{f}^{(k)}]\|^{2}+\beta^{2}m\sum_{i=1}^{m}\mathrm{Var}(\boldsymbol{\zeta}_{g_{i}}^{(k)})+\beta^{2}m\sum_{i=1}^{m}\|\mathbb{E}[\boldsymbol{\zeta}_{g_{i}}^{(k)}]\|^{2}\right)$
$\leq 2\alpha_{k}^{2}(\sigma_{1}^{2}+l_{f}^{2}+\beta^{2}m^{2}\sigma_{2}^{2}+\beta^{2}m^{2}l_{g}^{2})=\alpha_{k}^{2}L_{F}^{2}.$
Since $\mathrm{mod}(k,q)>0$, there exist nonnegative integers $t$ and $s<q$ such that $k=tq+s$. Summing Eq. 30 over $k=tq+1,\ldots,tq+s-1,tq+s$, and applying Eq. 29, we have
@@
The proof is then completed by combining Case (i) and Case (ii).
The following corollary is obtained immediately from Eq. 32 by noticing $r_{g}^{(tq)}=0$ if $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over $N$ data points and ${\mathcal{B}}_{tq}$ contains all the data points.
@@
Under 1 and 3, suppose that $\mathbb{P}_{{\mathbf{g}}}$ is a uniform distribution over a finite dataset $\Xi$ of size $N$. Let ${\mathcal{B}}_{k}=\Xi$ if $\mathrm{mod}(k,q)=0$. Then $r_{g}^{(k)}=0$ if $\mathrm{mod}(k,q)=0$ and $r_{g}^{(k)}\leq\sum_{i=1}^{q}\frac{\alpha_{\lfloor k/q\rfloor q+i-1}^{2}L_{F}^{2}L_{g}^{2}}{S_{2}}$ otherwise, where $L_{F}$ is given in Lemma 3.1.
Next, we bound the difference between the Moreau envelope at two consecutive iterates. Our analysis follows that of [20, Theorem 3.1], with the key modification in how we handle and bound the term $\mathbb{E}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{\mathbf{x}}^{(k)},\beta{\mathbf{w}}^{(k)}\right\rangle\right]$.
@@
Under 1 and 3, suppose $\{{\mathbf{x}}^{(k)}\}_{k=0}^{Kq-1}$ and $\{{\mathbf{u}}^{(k)}\}_{k=0}^{Kq-1}$ are generated by Algorithm 1. With $L_{F}$ given in Lemma 3.1, it holds
@@
$\leq$ $\mathbb{E}_{k}\left[F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\left\|\widehat{{\mathbf{x}}}^{(k)}-\left({\mathbf{x}}^{(k)}-\alpha_{k}\left(\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right)\right)\right\|^{2}\right]$
$=$ $F^{\nu}(\widehat{{\mathbf{x}}}^{(k)})+\rho\left\|{{\mathbf{x}}}^{(k)}-\widehat{{\mathbf{x}}}^{(k)}\right\|^{2}+2\rho\alpha_{k}\mathbb{E}_{k}\left[\left\langle\widehat{{\mathbf{x}}}^{(k)}-{{\mathbf{x}}}^{(k)},\boldsymbol{\zeta}_{F^{\nu}}^{(k)}+\beta{\mathbf{w}}^{(k)}\right\rangle\right]$
@@
where the first inequality follows directly from the definition of $F^{\nu}_{\frac{1}{2\rho}}$, and the last one uses $\mathbb{E}_{k}[\boldsymbol{\zeta}_{F^{\nu}}^{(k)}]\in\partial F^{\nu}({\mathbf{x}}^{(k)})$ by 3(b) and the $\rho$-weak convexity of $F^{\nu}$ by Lemma 2.7.